OpenStudy (anonymous):

What is the volume of a cone with a height of 6 centimeters and a radius of 2 centimeters (use 3.14 for pi )?

OpenStudy (anonymous):

i still dont understand /:

OpenStudy (anonymous):

so its: 6 x 4 x 3.14 . but i dont know what to do with the 1/3 .

OpenStudy (anonymous):

\[\LARGE V=\frac13 \cdot 3.14\cdot (2cm )^2\cdot 6cm\] \[\LARGE V=\frac13 \cdot 3.14\cdot 24 cm^3\] \[\LARGE V=25.12 cm^3\]

OpenStudy (anonymous):

multiply them, then divide... (use calculator)